OpenStudy (anonymous):

what is the answer to the following equation? 3x-4y=3 and 4y=x-9

OpenStudy (bahrom7893):

3x - 4y = 3 4y = x - 9 <-- You know that 4y is x - 9 plug that into the equation above

OpenStudy (bahrom7893):

3x - ( x - 9 ) = 3 3x - x + 9 = 3 2x + 9 = 3; 2x = 3 - 9; 2x = -6

OpenStudy (bahrom7893):

x = -6 / 2 = -3. Now you know that x = -3, so plug that into: 4y = x - 9 4y = -3 - 9 4y = -12 y = -3